​How many leaves does a full 3-ary tree with 100 vertices have?

How many leaves does a full 3-ary tree with 100 vertices have?
Answer #1

A full 3 - ary tree with 100 internal vertices has

L = ( 3 - 1 ) * 100 + 1

= 2 * 100 + 1

= 200 + 1

= 201 leaves.